Calculate the amount of theoretical raw material required for 1000 liters of extra neutral ethyl alcohol to be produced for use in acetic acid production.
Answer #1

Answer: Ethyl alcohol is prepared industrially by hydrating ethene in the presence of conc. H2SO4 which is as follows:

C_{2}H_{4}+H_{2}O\overset{conc. H_{2}SO_{4}}{\rightarrow} C_{2}H_{5}OH

Now, 1000 L of ethyl alcohol = 1000/22.4 = 44.64 moles of ethyl alcohol.

As it can be seen that 1 mole of ethyl alcohol theoretically requires one mole of ethene and one mole of water, so

Amount of ethene required = 44.64 moles = 44.64*22.4 = 1000 L

Amount of water required = 1000 L.